OpenStudy (anonymous):

The perimeter of a triangle is 61 centimeters. The longest side is 5 centimeters less than the sum of the other two sides. the shortest side is 14 centimeters less than the longest side. Find the lengths of each side of the triangle.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

a+b+c = 61 c = a+b -5 a = c-14 system of 3 equations solve using substitution

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

A=14 B=19 C=28

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

How do I rate your answer

OpenStudy (anonymous):

I appreciate the help really

OpenStudy (dumbcow):

click on blue "good answer" button to right of my name